OpenStudy (anonymous):

Please Help!!! D: Evaluate the Integral of (4u^2- 5u^-1)du

OpenStudy (anonymous):

use the standard integrals integral of a x^n = a x^(n + 1) / (n + 1) integral of x^-1 = ln x

OpenStudy (anonymous):

how would I write that out?

OpenStudy (anonymous):

in the case of 4 u^2 its [4 u^(2 + 1) ] / (2 + 1) simplify that

OpenStudy (anonymous):

so that would simplify to 4u^3/3... right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

now do the same to 5u^-1 which would simplify to 5?

OpenStudy (anonymous):

no - check my first pos

OpenStudy (anonymous):

the formula does not work for u^(-1)

OpenStudy (anonymous):

so 5 ln |u| ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

+c ?

OpenStudy (anonymous):

yes + constant of integration

OpenStudy (anonymous):

so my overall answer is 4u^3/3-5ln|u|+c ?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

let me see if the program that my online hw uses accepts it

OpenStudy (anonymous):

ok cool it looks like it took it! heres the next one...

OpenStudy (anonymous):

Evaluate the Integral of \[17\sqrt[3]{x}dx\]

OpenStudy (anonymous):

Thank you for the help with the last one btw!